OpenStudy (liliakarina):

A restaurant wants to determine the demand on cow meat, so that the orders per month can be made to the provider and frequency with which he must deliver. This board shows the results: Day 1 2 3 4 5 Meat kilo 29 33 25 28 What's the option that determines the relative frequency (fr) as proportion? a-1 2 2 3 4 b-3.33 6.66 6.66 10.00 13.33 c-0.033 0.066 0.066 0.1 0.133 d-0.01 0.02 0.02 03 0.04

OpenStudy (anonymous):

Hi .. I'm not quite sure that you are give us the right choices but I try to solve it : Total meat in kilo = 29 + 33 + 25 + 28 =115 \[\frac {29}{115},\frac {33}{115},\frac {25}{115},\frac {28}{115}=0.243,0.287,0.217,.243\] if we add all Fr then when we must have 1 also you can check this if you like : http://www.referenceforbusiness.com/management/Sc-Str/Statistics.html#b regards